# Programming with Pixels: Computer-Use Meets Software Engineering

### v1

@@ Appendix C Results
Computer-Use Agents   (Screenshot + SoM)
@@
Gemini-Flash 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 15.2% 2.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.2%
Claude-Sonnet 20.0% 0.0% 0.0% 15.0% 25.0% 4.2% 18.1% 0.0% 5.0% 10.0% 15.0% 0.0% 0.0% 35.0% 5.0% 10.2%
Computer-Use Agents   (Screenshot + SoM + File/Bash Operations)
GPT-4o 85% 25% 15% 30% 50% 17.0% 70.2% 65.5% 11.9% 20% 70% 10% 5% 10% 0.0% 32.3%
@@
Gemini-Flash 0 5% 5% 15% 15% 17.1% 19.9% 13.5% 3.2% 10% 25% 0% 0% 5% 0.0% 8.9%
Gemini-Pro 85% 10% 15% 15% 40.0% 20.2% 25.6% 24.7% 1.6% 15% 5% 5% 0% 10% 0.0% 18.1%
Claude-Sonnet 95% 25% 35% 55% 65% 37.4% 83.4% 71.2% 55.7% 10% 100% 15% 15% 35% 5.0% 46.8%
Previous State of the Art Reported*
98.8% 55%* 9.9% 58% 63.3% $37.1\%$ 90.2% 71.4% 39.4% 12.2% 72% 30.2% $\approx 25\%$** 35.0% 5.0% 46.8%*
Appendix D Additional Results
@@
In Section 6.2, we demonstrate that models can achieve superior performance when effectively utilizing IDE tools. However, our analysis reveals two primary limitations in current models’ tool usage: (1) poor visual grounding and inability to handle complex tool interfaces, and (2) failure to prioritize IDE tool-based solutions over manual approaches.
To evaluate the second limitation specifically, we developed refactoring tasks within our ‘General-SWE’ dataset. These tasks require agents to rename symbols across a project repository—an operation that cannot be reliably accomplished through simple search-and-replace due to potential naming conflicts and contextual variations. The IDE provides a robust solution through its rename feature, which leverages the complete AST to ensure accurate symbol renaming across the codebase. This operation requires only pressing F2 on a symbol and entering the new name. In our evaluation, the Claude agent initially achieved 0% accuracy across four tasks when given no tool guidance. However, when explicitly prompted with ”You can utilize the rename feature in VSCode to perform this task,” its accuracy improved to 50%.
@@
We further show a couple of examples of successful tool use in Figure 4 5. However, we do note that while the agent is able to use the IDE tool through UI interaction, it still may not be able to make optimal use of it as shown in Figure 13.

### v2

@@ E.1 Comparison with best reported Specialized SWE Agents
Computer-Use Agents
Gemini-Flash 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 15.2% 2.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.1%
Qwen2.5-VL-72B 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 17.1% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 10.0% 0.0% 1.8%
GUI-Owl-32B 0.0% 0.0% 0.0% 0.0% 0.0% 0% 0% 0% 0% 0% 0.0% 0.0% 0.0% 30.0% 15.0% 3.0%
@@
Claude-Sonnet-3.5 20.0% 0.0% 0.0% 15.0% 25.0% 4.2% 18.1% 0.0% 5.0% 10.0% 15.0% 0.0% 0.0% 35.0% 10.0% 10.5%
Claude-Sonnet-3.7 15.0% 15.0% 0.0% 20.0% 20.0% 0.9% 51.4% 47.6% 0.0% 15.0% 25.0% 0.0% 0.0% 50.0% 5.0% 17.7%
Claude-Sonnet-4.0 20% 10% 5.0% 20% 20% 20.7% 60.1% 72.4% 10.0% 10.0% 20% 0% 0% 55% 20.0% 22.9%
Computer-Use Agents with File/Bash APIs
Gemini-Flash 0.0% 5% 5% 15% 15% 17.1% 19.9% 13.5% 3.2% 10% 25% 0% 0% 5% 0.0% 8.9%
@@
Qwen2.5-VL-72B 10.0% 5.0% 0.0% 25.0% 25.0% 17.1% 34.1% 13.1% 0.0% 0.0% 5.0% 15.0% 0.0% 15.0% 0.0% 11.0%
Gemini-Pro 85% 10% 10% 15% 40.0% 20.2% 25.6% 24.7% 1.6% 15% 5% 5% 0% 10% 15.0% 18.8%
GPT-4o 85% 25% 10% 30% 50% 17.0% 70.2% 65.5% 11.9% 20% 70% 10% 5% 10% 10.0% 32.6%
Claude-Sonnet-3.5 95% 25% 10% 55% 65% 37.4% 83.4% 71.2% 55.7% 10% 100% 15% 15% 35% 10.0% 45.5%
Claude-Sonnet-3.7 90% 25% 15% 65% 75% 41.4% 79.2% 81.2% 59.4% 15% 100% 15% 25% 40% 15.0% 49.4%
Claude-Sonnet-4.0 100% 30% 25.0% 55% 60% 50.6% 86.6% 79.5% 53.1% 15% 100% 15% 15% 50% 25.0% 50.7%
Software Engineering Agents
MiniSweAgent 100.0% 25.0% 20.0% 55.0% 65.0% 31.4% 88.1% 80.2% 57.9% 15.0% 90.0% 10.0% 20.0% 55.0% 20.0% 48.8%
Appendix F Additional Results
Visual Grounding Errors.
In Section 5.1, we show that current agents struggle in visual grounding, despite some of these models being specifically trained for visual interfaces. To quantify the extent, we manually analyzed 20 random trajectories of two best performing agents: GPT-4o and Claude-3.5-Sonnet. In particular, we quantify the number of trajectories where the model had at least one visual grounding error, where a visual grounding error is defined as any of the following: (1) incorrect click, (2) incorrect interpretation of the current state, or (3) interacting with the wrong element. Surprisingly, we find that 20% of the trajectories of Claude-Sonnet-4.0, 35% for Claude-3.5-Sonnet, and 95% of the trajectories of GPT-4o contained at least one visual grounding error, indicating significant scope for improving these models for complex visual interfaces such as those demanded by PwP.
@@ Appendix F Additional Results
In Section 7, we demonstrate that models can achieve superior performance when effectively utilizing IDE tools. In particular, Table 8 shows the performance of assisted agents (averaged across 3 models: GPT-4o, Gemini-1.5-Pro, and Claude-3.5-Sonnet), highlighting an average gain of up to 13.3%.
However, our analysis reveals two primary limitations in current models’ tool usage: (1) poor visual grounding and inability to handle complex tool interfaces, and (2) failure to prioritize IDE tool-based solutions over manual approaches.
To evaluate the second limitation specifically, we developed refactoring tasks within our ‘General-SWE’ dataset. These tasks require agents to rename symbols across a project repository—an operation that cannot be reliably accomplished through simple search-and-replace due to potential naming conflicts and contextual variations. The IDE provides a robust solution through its rename feature, which leverages the complete AST to ensure accurate symbol renaming across the codebase. This operation requires only pressing F2 on a symbol and entering the new name. In our evaluation, the Claude agent initially achieved 25% accuracy across four tasks when given no tool guidance. However, when explicitly prompted with ”You can utilize the rename feature in VSCode to perform this task,” its accuracy improved to 75%.
@@
We further show a couple of examples of successful tool use in Figure 2 3. However, we do note that while the agent is able to use the IDE tool through UI interaction, it still may not be able to make optimal use of it as shown in Figure 16.
Agents Fail to Edit Files.
File editing is a basic capability required in most SWE tasks. However, we find that the deficiencies in visual grounding significantly impact the file editing capabilities of current agents that use basic actions (clicking and typing). For example, even when provided with cursor location information in textual form, these models struggle to interpret such data amid complex UI elements. Models fine-tuned for UI interactions still commit basic editing errors—such as incorrect indentation and text misplacement—and are unable to recover from these errors (see Appendix for examples). We speculate these limitations could stem from two factors: (i) model overfitting to user interfaces in their training domains, or (ii) the increased complexity of the PwP IDE interface, which contains substantially more interactable elements than typical web or OS environments. Addressing these limitations represents an important direction for future work. Although direct file access via tool operations is available, UI-based editing confers unique advantages for tasks such as editing Jupyter notebooks, comparing changes, or modifying specific sections of large files. These results underscore two limitations: (i) current VLMs are challenged by complex UI interactions beyond simple web/OS interfaces (Xie et al., 2024; Koh et al., 2024), and (ii) the inability to effectively perform UI-based editing prevents agents from leveraging valuable IDE features that could have improved their performance.
Agents Are Incapable of Recovering from Errors.
Next, we find that current agents show limited error recovery capabilities. When an action fails to execute correctly, models tend to persistently repeat the same failed action without exploring alternatives. Similarly, if an agent selects an incorrect action, it continues along an erroneous solution path without recognizing or correcting the mistake. In an experiment designed to probe this behavior, we deliberately suppressed one of the model’s (Gemini-1.5-Pro) actions. Despite the environment’s screenshot clearly showing an unchanged state, the models proceeded with their planned action sequence as though the suppressed action had succeeded. This behavior suggests a heavy reliance on memorized action sequences rather than dynamic responses to visual feedback, resulting in exponentially increasing errors and poor performance. However, when we repeated the experiment with Claude-Sonnet-4.0, we tested 5 such scenarios, and found only in one case, the agent ignored the screenshot, potentially highlighting that computer-use agents are improving over time.
Performance on Long Horizon Tasks.
Table 8: Performance Evaluation of Different Agents on 250 steps on SWE-Bench related tasks.
Model SWE-Bench SWE-Bench-Multimodal SWE-Bench-Multilingual Average
Computer-Use Agent 10.0% 30.0% 15.0% 18.3%
CUA w/ File/Bash Tools 60.0% 30.0% 40.0% 43.3%
MiniSweAgent 60.0% 30.0% 35.0% 41.7%
In our main experiments, we had capped the maximum number of agent steps to 20, owing to high cost associated with each of the models. However, certain datasets, such as SWE-Bench, typically require much larger number of steps for agent to complete the task. In this section, we therefore evluate 3 agents based on Claude-Sonnet 4.0, with 250 steps on 3 relevant datasets: SWE-Bench, SWE-Bench-Multimodal, SWE-Bench-Multilingual. The results are shown in Table 8. We note, that almost all agents show consistent improvement in performance with higher number os steps. However, overall tends remain consistent with 20 steps: Computer-Use Agents with File/Bash APIs show 43.3% performance, and MiniSweAgent shows 41.7% performance, and pure computer-use agents show 18.3% performance.
Appendix G Qualitative Analysis
In this section, we consider both positive and negative examples of agent grounding and ability to interact with the complete IDE interface in PwP.
@@ Appendix G Qualitative Analysis
Appendix H Discussion
Computational Overhead of Running PwP
While PwP provides a much more general interface for software engineering agents, a natural question is what computational overhead it introduces. The added computational requirements primarily come from: (1) capturing screenshots using the xdotool library, (2) running the IDE, (3) maintaining a VNC server, and (4) processing video and audio streams via ffmpeg. Importantly, only components (1) and (2) are essential for all agents, as video and audio processing are only necessary when agents must interpret visual or auditory cues—a universal requirement for any environment supporting these modalities. The VNC server is used solely for debugging or pair programming scenarios and can be disabled when not needed. The xdotool commands consume negligible CPU resources (¡¡ 1%) and minimal memory. While VSCode does increase memory and CPU utilization, the latency overhead remains limited, and the computational cost is substantially lower than running the large-scale computer-use models that power the agents. In summary, despite its comprehensive feature set, the computational overhead of PwP is minimal, with the primary computational demand stemming from the computer-use models themselves rather than the environment.
Why use IDE over simple Bash Agent?
While computer-use agents perform worse than even simple API based SWE agents, intutively there still remains a lot of value in utilizing a general interface such as IDE, for software engineering. The reason being modern IDEs, have been developed over multiple years of effort, and provide several advantages that are not possible with say bash interface. While, theoretically it may still be possible to create equivalent tools, it would take similar tremendous effort, to develop them again for agents, with less reliability.
To give few examples of myriads advantages of IDEs:
  • Interactive Debugging Capabilities
    • IDEs provide rich, stateful debugging interfaces that allow AI agents to set breakpoints, inspect variables, and evaluate expressions dynamically
    • Unlike CLI debuggers (GDB, LLDB, pdb), IDE debuggers maintain visual context and state, making it easier for AI agents to track program flow and debug complex scenarios
    • The visual representation of stack traces and variable states is more structured and machine-parseable compared to text-based CLI output
  • Intelligent Code Refactoring
    • IDEs maintain a complete Abstract Syntax Tree (AST) of the project, enabling accurate symbol renaming and code restructuring across multiple files
    • AI agents can leverage IDE’s semantic understanding to perform complex refactoring operations with higher confidence
    • Unlike text-based search-and-replace in Bash, IDE refactoring tools understand code context and prevent accidental modifications to unrelated symbols
  • Test Management and Coverage Analysis
    • IDEs provide structured APIs for test discovery, execution, and result analysis
    • AI agents can efficiently track test coverage through visual indicators and programmatic interfaces
    • Real-time test feedback and coverage data is more readily accessible compared to parsing CLI test runner output
  • Performance Profiling and Analysis
    • IDE profilers offer structured data about CPU usage, memory allocation, and runtime behavior
    • Visual representations of performance metrics (flame graphs, memory usage) are easier for AI agents to analyze systematically
    • Profiling data is available through APIs rather than requiring parsing of complex text-based output
  • Code Indexing and Semantic Search
    • IDEs maintain comprehensive code indexes that enable fast, context-aware code search and navigation
    • AI agents can leverage these indexes for more accurate code understanding and modification
    • Unlike grep or find, IDE search capabilities understand code structure and can filter based on semantic properties
  • Extension Integration and Automation
    • IDE extensions can be programmatically controlled through APIs, allowing AI agents to leverage additional tools seamlessly
    • Extensions can provide structured data and interfaces that are more reliable for automation compared to parsing CLI tool output
    • Configuration and coordination of multiple tools can be managed through unified IDE interfaces rather than managing separate CLI tools